# View Selection for 3D Captioning via Diffusion Ranking

### v1

@@ Abstract
Scalable annotation approaches are crucial for constructing extensive 3D-text datasets, facilitating a broader range of applications. However, existing methods sometimes lead to the generation of hallucinated captions, compromising caption quality. This paper explores the issue of hallucination in 3D object captioning, with a focus on Cap3D Luo et al. (2023a) method, which renders 3D objects into 2D views for captioning using pre-trained models. We pinpoint a major challenge: certain rendered views of 3D objects are atypical, deviating from the training data of standard image captioning models and causing hallucinations. To tackle this, we present DiffuRank, a method that leverages a pre-trained text-to-3D model to assess the alignment between 3D objects and their 2D rendered views, where the view with high alignment closely represent the object’s characteristics. By ranking all rendered views and feeding the top-ranked ones into GPT4-Vision, we enhance the accuracy and detail of captions, enabling the correction of 200k captions in the Cap3D dataset and extending it to 1 million captions across Objaverse and Objaverse-XL datasets. Additionally, we showcase the adaptability of DiffuRank by applying it to pre-trained text-to-image models for a Visual Question Answering task, where it outperforms the CLIP model.
@@ 1 Introduction
Recent advancements in generative models have shown remarkable performance in both image Saharia et al. (2022); Betker et al. (2023) and video Brooks et al. (2024) domains, driven by the availability of extensive captioned datasets. Despite these successes, extending generative modeling to 3D domains has been challenging due to the scarcity of high-quality 3D-text pairs. This gap has been partially bridged by Cap3D Luo et al. (2023a), which generates captions for 3D objects by rendering them into 2D images and employing image-based captioning models, further refined by Large Language Models (LLMs) to synthesize captions. Cap3D has contributed 660k captions for the Objaverse dataset Deitke et al. (2023a), facilitating developments in Text-to-3D methods Yariv et al. (2023); Li et al. (2023a), Image-to-3D methods Xu et al. (2023a); Zhao et al. (2023a), robotic simulator Wang et al. (2023a) and learning Qi et al. (2024), and the pre-training of 3D LLMs Xu et al. (2023b); Zhou et al. (2023); Panagopoulou et al. (2023).
Despite the utility of Cap3D, our analysis reveals that a significant portion of Cap3D captions includes inaccurate and hallucinated information, potentially compromising model training Tang et al. (2023a). Upon inspection, we found that the key is the rendered view: as Cap3D adheres to the Objaverse’s default orientation for 3D objects, it positions the rendering cameras horizontally based on heuristic hyperparameters. Some of the renderings are hard to distinguish even for humans, which cannot be handled by the existing captioning models Li et al. (2023b). Consequently, when these challenging views are included, even advanced captioning models like GPT4-Vision Achiam et al. (2023) may generate erroneous information, as illustrated in the second and fourth rows of Figure 1.
@@
To address this, we introduce DiffuRank, an approach for ranking rendered views based on 3D priors learned by pre-trained diffusion models. By leveraging a pre-trained text-to-3D diffusion model Jun and Nichol (2023), DiffuRank evaluates the alignment between the captions of each view and the information of its corresponding 3D object. The underlying premise is that captions generated from rendered views that closely match the object’s 3D information will exhibit a higher alignment, suggesting these views are more representative of the object. Consequently, DiffuRank promotes that the preferable views (Figure 2) for captioning are those that better reflect the true essence of the 3D objects, leading to more accurate and truthful captions.
Specifically, we first employ image-based captioning models to caption all candidate rendered views, and then perform multiple iterations over diffusion model objective to obtain average score estimation for all captions conditional on the same 3D object feature, Gaussian noise, and timestamps. This score gauges the alignment between the captions and the corresponding 3D object feature. Following this, we rank the views based on their scores and forward the top-N rendered views to GPT4-Vision for the final caption generation. Our evaluations through human studies indicate that captions produced with DiffuRank, in conjunction with GPT4-Vision, are of significantly higher quality and exhibit fewer inaccuracies compared to those generated by Cap3D. Moreover, our captions are usually richer in detail and fewer hallucinations when using only 6 rendered views than those produced using GPT4-Vision alone across all 28 rendered views or views selected based on default object orientations.
@@
Additionally, we extend DiffuRank to the 2D domain, demonstrating its effectiveness in the challenging Visual Question Answering task Tong et al. (2024) when combined with a text-to-2D diffusion models Rombach et al. (2022), surpassing the zero-shot performance of CLIP Radford et al. (2021).
@@
We extend the Cap3D caption dataset Luo et al. (2023a) from 660k to 1M across the whole Objaverse Deitke et al. (2023a) and subset of Objavere-XL highquality set Deitke et al. (2023b). The captions are complemented with point clouds and rendered images, including camera, depth, and MatAlpha details, all releasing under the ODC-By 1.0 license.
We proposed DiffuRank which shown ability to model the alignment between 3D object and its 2D rendered views via a pre-trained Text-to-3D model and a captioning model. Additionally, we extend DiffuRank to 2D domain, and demonstrate DiffuRank beats CLIP on the VQA task Tong et al. (2024) with the help of a pre-trained text-to-2D diffusion model Rombach et al. (2022).
@@ 2 Related Work
Recent advancements introduced by Objaverse have significantly enriched the field of 3D object research. By integrating a comprehensive set of 3D objects with descriptive captions from Cap3D, a wide array of 3D applications has been enabled. These include Text-to-3D methods Yariv et al. (2023); Li et al. (2023a); He et al. (2023); Li et al. (2023c); Mercier et al. (2024), Image-to-3D conversion techniques Xu et al. (2023a); Zhao et al. (2023a), enhancements in robot learning Wang et al. (2023a); Qi et al. (2024), the pre-training of 3D language models Xu et al. (2023b); Zhou et al. (2023); Qi et al. (2023); Liu et al. (2024a); Chen et al. (2023a), and the development of language models capable of processing diverse modalities Han et al. (2023); Panagopoulou et al. (2023); Chen et al. (2024).
Despite these advancements, we identified issues with hallucination contents in the captions provided by Cap3D. This discovery aligns with findings from concurrent research Tang et al. (2023a); Liu et al. (2023a); Kabra et al. (2023), pinpointing inaccuracies in Cap3D captions. Our investigation reveals that the root cause of these inaccuracies is attributed to atypical rendered views, which lead to failures in captioning models. These failures are exacerbated as text summarization models (GPT4) are unable to rectify these errors. To address this challenge, we introduce DiffuRank that selects rendered views capturing the essential characteristics of 3D objects. Furthermore, we utilize the recent advancements in vision-language models, specifically GPT4-Vision, to provide holistic captions for 3D objects. We release our dataset under ODC-By 1.0 license to enable research and commercial usage, and hope facilitate related 3D-Text research Jain et al. (2021); Poole et al. (2022); Lin et al. (2022); Sanghi et al. (2022); Zhu and Zhuang (2023); Wang et al. (2023b); Chen et al. (2023b); Lorraine et al. (2023); Li et al. (2023a); Yi et al. (2023); Li et al. (2023c); Luo et al. (2023b); Ding et al. (2023); Chen et al. (2023c); Michel et al. (2022); Wei et al. (2023); Chen et al. (2023d); Nichol et al. (2022); Liu et al. (2023b, c); Melas-Kyriazi et al. (2023); Tang et al. (2023b); Shi et al. (2023); Xu et al. (2023a); Chen et al. (2023e); Shi et al. (2023).
@@ 2.2 Diffusion Model
Our proposed DiffuRank leverages denoising diffusion objective Sohl-Dickstein et al. (2015); Song and Ermon (2019); Ho et al. (2020) to model the alignment between the input and output modalities. By using pre-trained text-to-3D Nichol et al. (2022); Jun and Nichol (2023) and text-to-2D Saharia et al. (2022); Betker et al. (2023); Peebles and Xie (2023) diffusion models, we can model the alignment between given 3D object/image for a set of possible captions (text descriptions) as detailed in Section 3.2. In our listed algorithm 1, we adopt the objects as used in Shap·E Jun and Nichol (2023), where is data sampled from data distribution , is Gaussian noise, and t is timestamp. We also adopt the alternative but equivalent objective, , when we adopt the text-to-2D model, stable-diffusion, in Section 5.3.
DiffuRank is related to score sampling distillation proposed in Poole et al. (2022), while we do not compute gradients but sampling loss to accumulate scores estimation for ranking. Our findings also relate to works which leverage pre-trained diffusion models for downstream tasks, such as image classification Mukhopadhyay et al. (2023); Li et al. (2023d), semantic segmentation Zhao et al. (2023b), visual grounding Liu et al. (2023d), depth prediction Saxena et al. (2023); Zhang et al. (2024), and other low-level computer vision tasks Du et al. (2023).
When applying our method to the 2D domain, we discovered that our algorithm aligns closely with the insights of the approach presented in Li et al. (2023d). Consequently, our method can be considered an expansion of the findings from Li et al. (2023d), extending its applicability from 2D classification to broader domains and tasks, including the use of a pre-trained text-to-3D diffusion model and a captioning model to estimate the alignment between 3D objects and their 2D rendered views, as well as the application of a pre-trained text-to-2D diffusion model to solve Visual Question Answering (VQA) tasks. Note that DiffuRank, which requires extensive sampling for each candidate as outlined in step 3 of Algorithm 1, may not be suitable for tasks with numerous options. More discussions are included in Section 6.
@@ 3 Method
In this section, we analyze the issues with atypical rendered views leading to hallucinations in Cap3D captions, motivating our proposed DiffuRank, a approach for selecting informative rendered views with 3D priors learned from a diffusion model. We then detail DiffuRank’s formulation and describe our novel 3D captioning framework that integrates GPT4-Vision.
@@ 3.1 Issues in Cap3D
Firstly, we revisit the Cap3D pipeline, which unfolds across four stages. Initially, it renders a set of 2D views for each 3D object. Subsequently, image captioning is applied to generate preliminary descriptions (5 captions for each image). Then, the CLIP model is utilized in the third stage for selecting the best aligned caption for each image to filter out inaccuracies. The process culminates with an LLM synthesizing captions from various perspectives into a unified comprehensive caption.
However, the captioning of rendered views (the combined second and third stages) for given 3D objects can falter with atypical views, producing captions that diverge significantly from the actual 3D object. In the worst-case scenarios, each rendering view might correspond to an incorrect object, leading to compounded errors when these captions are summarized by GPT4. One example is shown in Figure 3. Since GPT4 operates solely on text, it cannot correct these inaccuracies, resulting in captions riddled with hallucinated details.
Addressing this challenge is non-trivial, as determining the appropriate view for any given 3D object is complex. While measuring the geometric properties of 3D objects and computing their principal direction is feasible, positioning the camera orthogonally, as shown in the bottom-left example of Figure 2, is often suboptimal. Hence, we propose DiffuRank, which learns 3D priors from data to filter informative rendered views by leveraging a pre-trained text-to-3D model. Our experiments demonstrate that DiffuRank efficiently enhances caption quality and reduces hallucinations with fewer renderings than using all available views.
@@ 3.2 DiffuRank Formulation
DiffuRank leverages a pre-trained text-to-3D diffusion model to rank rendered views based on their alignment with both captions and the corresponding 3D information.
For a given 3D object , assuming a set of candidate captions and the pre-trained model , the training objective of this pre-trained diffusion model is predicting a 3D object based on a text description , i.e., modeling the score function of the data distribution . Specially, the diffusion model aims to minimize
based on a text description , where the noised input , for timestamp , and randomly sampled Gaussian noises , with being a hyper-parameters defined by the noise schedule Ho et al. (2020). Our tuition here is simple: a caption closely aligned with the given 3D object in terms of characteristics (e.g. structure, colors, textures, etc), should aid the diffusion model in making accurate predictions starting from the same noised input , resulting in a lower score matching loss. By sampling multiple sets of for the same set of captions , we can measure the alignment between the 3D object and captions via the average loss.
Initially, we generate candidate captions for by rendering it into multiple views and generating captions with a captioning model . This captioning procedure aims to maximize the joint likelihood of the model distribution over the image and generated captions . Thus, we estimate the alignment between the 3D object and all captions of the same rendering , which is proportional to . Then, we write down the whole pipeline in Algorithm 1.
Specifically, we adopted shap-E as the text-to-3D diffusion model in our paper, and the above should be , where is the encoder (transmitter in Jun and Nichol (2023)) to extract feature embeddings from given 3D object.
Furthermore, DiffuRank’s application is not confined to 3D captioning; because it is a general framework for measuring the alignment between two modalities received and output by a diffusion model. It can be seamlessly extended to other domains, such as 2D images. In section 5.3, we show an example where we apply DiffuRank to perform 2D VQA and beat CLIP model Radford et al. (2021).
@@ 3.3 New 3D Captioning Framework
With the proposed DiffuRank, we establish a new 3D captioning pipeline, as shown in Figure 3. For given 3D object, we render it into 28 images, which are then captioned into 5 descriptions using an image-based captioning model. Following captioning, DiffuRank ranks the rendered views using a pre-trained text-to-3D model. This ranking enables the selection of the Top-6 rendered views for processing by a vision-language model, resulting in holistic captions that describe structure, form, color, texture, and more, with enhanced accuracy and detail.
To elaborate, our methodology integrates two distinct rendering strategies, as illustrated in Figure 4. The first strategy, derived from Cap3D Luo et al. (2023a), renders objects into 8 views against a uniform grey background, arranged horizontally around the object’s default orientation, with Blender ray-tracing render engine ‘CYCLES’. Concurrently, we apply a second technique from Shap·E Jun and Nichol (2023), where 20 views are generated through randomized sampling after object normalization, set against a transparent background, with Blender real-time engine ‘EEVEE’. These 20 views, created following the Shap·E methodology, are instrumental in forming Shap·E latent codes, i.e. in Section 3.2. Altogether, this approach results in 28 distinct views for each object. Additionally, as grey and transparent backgrounds may accentuate or obscure details variably across objects, we observed that DiffuRank adeptly selects the views with the proper background that most effectively highlight object features, without manual intervention. Some examples are included in Appendix B.
Following this, the captioning model, BLIP2 Li et al. (2023b), is employed to generate five captions for each view. These captions, alongside the pre-trained text-to-3D diffusion model, Shap·E Jun and Nichol (2023), and the previously derived 3D latent code , undergo analysis in the DiffuRank process, as detailed in Algorithm 1. Subsequent to DiffuRank, the six views that demonstrate the highest alignment scores are chosen to input into GPT4-Vision for caption generation.

### v2

# https://huggingface.co/datasets/tiange/Cap3D View Selection for 3D Captioning via Diffusion Ranking

@@ Abstract
Scalable annotation approaches are crucial for constructing extensive 3D-text datasets, facilitating a broader range of applications. However, existing methods sometimes lead to the generation of hallucinated captions, compromising caption quality. This paper explores the issue of hallucination in 3D object captioning, with a focus on Cap3D [36] method, which renders 3D objects into 2D views for captioning using pre-trained models. We pinpoint a major challenge: certain rendered views of 3D objects are accidental [16], deviating from the training data of standard image captioning models and causing hallucinations. To tackle this, we present DiffuRank, a method that leverages a pre-trained text-to-3D model to assess the alignment between 3D objects and their 2D rendered views, where the view with high alignment closely represent the object’s characteristics. By ranking all rendered views and feeding the top-ranked ones into GPT4-Vision, we enhance the accuracy and detail of captions, enabling the correction of 200k captions in the Cap3D dataset and extending it to 1.5 million captions across the entire Objaverse dataset and a portion of the Objaverse-XL high-quality subset. Additionally, our dataset includes point clouds with 16,384 colored points and 20 rendered images per caption, along with intrinsic and extrinsic camera parameters, depth data, and masks. Beyond datasets, we showcase the adaptability of DiffuRank by applying it to pre-trained text-to-image models for visual question answering tasks, where it outperforms the CLIP model.
@@ 1 Introduction
Recent advancements in generative models have shown remarkable performance in both image [51, 2] and video [4] domains, driven by the availability of extensive captioned datasets. Despite these successes, extending generative modeling to 3D domains has been challenging due to the scarcity of high-quality 3D-text pairs. This gap has been partially bridged by Cap3D [36], which generates captions for 3D objects by rendering them into 2D images and employing image-based captioning models, further refined by Large Language Models (LLMs) to synthesize captions. Cap3D has contributed 660k captions for the Objaverse dataset [13], facilitating developments in Text-to-3D [66, 24], Image-to-3D [65, 69], robot simulator [62] and learning [47], and the pre-training of 3D LLMs [64, 71, 44].
Despite the utility of Cap3D, our analysis reveals that a significant portion of Cap3D captions contain inaccurate or hallucinated details, potentially compromising model training [58]. Upon inspection, we identified one of the root causes as accidental views – special camera positions where coincidental feature alignments make the projection uninformative or misleading for recognition. Examples include the first “baseball" image, which actually depicts a green creature, and the third “cube" image, which is in fact an indoor design, both shown in Figure 1. Because Cap3D inherits Objaverse’s default object orientation, its cameras sit on a fixed horizontal ring, a heuristic that inevitably samples such accidental views. These accidental renderings are hard even for humans, and current captioners [25] struggle to describe them faithfully. Consequently, when these challenging views are included, even advanced captioning models like GPT4-Vision [1] may generate erroneous information, as illustrated in Figure 1.
@@
To address this, we introduce DiffuRank, an approach for ranking rendered views with pre-trained diffusion models. By leveraging a pre-trained text-to-3D diffusion model [21], DiffuRank evaluates the alignment between the captions of each view and the corresponding 3D object’s information. The underlying premise is that captions generated from rendered views that closely match the object’s 3D information will exhibit a higher alignment, indicating these views are more representative of the object. Consequently, DiffuRank promotes the preferable views (Figure 2) for captioning as those that better reflect the true essence of the 3D objects, leading to more accurate and truthful captions.
Specifically, we first employ image-based captioning models to caption all candidate rendered views. We then utilize multiple iterations of a text-to-3D diffusion model objective to estimate an average alignment score, using each caption as a conditioning input for the same underlying 3D object feature, Gaussian noise, and timestamps. The intuition here is that captions better aligned with the 3D object feature provide more effective guidance during the diffusion process, resulting in higher alignment scores. This score effectively quantifies how well each caption aligns with the corresponding 3D object feature. Since the captions are generated from rendered views via captioning models trained through maximum likelihood estimation, the resulting alignment scores inherently reflect the alignment between the rendered views and the original 3D object. Higher scores indicate stronger alignment, implying fewer possibilities of being accidental views.
Following this scoring procedure, we rank the views based on their alignment scores and select the top-N rendered views for further caption generation using GPT4-Vision. Human evaluations demonstrate that captions produced using DiffuRank combined with GPT4-Vision exhibit significantly higher quality and fewer inaccuracies compared to captions generated by Cap3D. Additionally, our method consistently yields richer, more detailed captions and fewer hallucinations when employing just 6 strategically selected rendered views, as opposed to using GPT4-Vision directly on all 28 views or those selected by default object orientations.
@@
Further, we extend DiffuRank to the 2D domain, demonstrating its effectiveness in the challenging Visual Question Answering task [60] when combined with text-to-2D diffusion models [50], and surpassing the zero-shot performance of CLIP [49].
@@
We propose DiffuRank to address the issue of accidental views in 3D object renderings. In principle, DiffuRank estimates the alignment score between different modalities. Therefore, we extend DiffuRank to the 2D domain and demonstrate that, with the assistance of a pre-trained text-to-2D diffusion model [50], DiffuRank outperforms CLIP on the VQA task [60].
@@
We extend the Cap3D caption dataset [36] from 660k to 1.5M across the entire Objaverse [13] and a portion of the Objavere-XL high-quality subset [12]. Each caption is complemented with point clouds containing 16,384 colorful points and 20 rendered images, including camera, depth, and MatAlpha details. All data is released under the ODC-By 1.0 license and is available at https://huggingface.co/datasets/tiange/Cap3D.
@@ 2 Related Work
Recent advancements introduced by Objaverse have significantly enriched the field of 3D object research. By integrating a comprehensive set of 3D objects with descriptive captions from Cap3D, a wide array of 3D applications has been enabled. These include Text-to-3D methods [66, 24, 18, 26, 38], Image-to-3D conversion techniques [65, 69], enhancements in robot learning [62, 47], the pre-training of 3D language models [64, 71, 48, 28, 8], and the development of language models capable of processing diverse modalities [17, 44, 5].
Despite these advancements, we identified issues with hallucination contents in the captions provided by Cap3D. This discovery aligns with findings from concurrent research [58, 33, 22], pinpointing inaccuracies in Cap3D captions. Our investigation reveals that the root cause of these inaccuracies is attributed to accidental rendered views, which lead to failures in captioning models. These failures are exacerbated as text summarization models (GPT4) are unable to rectify these errors. To address this challenge, we introduce DiffuRank that selects rendered views capturing the essential characteristics of 3D objects. Furthermore, we utilize the recent advancements in vision-language models, specifically GPT4-Vision, to provide holistic captions for 3D objects. We release our dataset under ODC-By 1.0 license to enable research and commercial usage, and hope facilitate related 3D-Text research [20, 46, 27, 52, 73, 61, 7, 34, 24, 67, 26, 35, 14, 10, 39, 63, 6, 42, 31, 30, 37, 57, 54, 65, 9, 54].
@@ 2.2 Diffusion Model
Our proposed DiffuRank leverages denoising diffusion objective [55, 56, 19] to model the alignment between the input and output modalities. By using pre-trained text-to-3D [42, 21] and text-to-2D [51, 2, 45] diffusion models, we can model the alignment between given 3D object/image for a set of possible captions (text descriptions) as detailed in Section 3.2. In our listed algorithm 1, we adopt the objects as used in Shap·E [21], where is data sampled from data distribution , is Gaussian noise, and t is timestamp. We also adopt the alternative but equivalent objective, , when we adopt the text-to-2D model, stable-diffusion, in Section 5.3.
DiffuRank is related to score sampling distillation proposed in [46], while we do not compute gradients but sampling loss to accumulate scores estimation for ranking. Our findings also relate to works which leverage pre-trained diffusion models for downstream tasks, such as image classification [40, 23], semantic segmentation [70], visual grounding [32], depth prediction [53, 68], and other low-level computer vision tasks [15].
When applying our method to the 2D domain, we discovered that our algorithm aligns closely with the insights of the approach presented in [23]. Consequently, our method can be considered an expansion of the findings from [23], extending its applicability from 2D classification to broader domains and tasks, including the use of a pre-trained text-to-3D diffusion model and a 2D-image-based captioning model to estimate the alignment between 3D objects and their 2D rendered views, as well as the application of a pre-trained text-to-2D diffusion model to solve Visual Question Answering tasks.
@@ 3 Method
In this section, we analyze the issues with accidental rendered views leading to hallucinations in Cap3D captions, motivating our proposed DiffuRank, a approach for selecting informative rendered views with 3D priors learned from a diffusion model. We then detail DiffuRank’s formulation and describe our novel 3D captioning framework that integrates GPT4-Vision.
@@ 3.1 Issues in Cap3D
Firstly, we revisit the Cap3D pipeline, which unfolds across four stages. Initially, it renders a set of 2D views for each 3D object. Subsequently, image captioning is applied to generate preliminary descriptions (5 captions for each image). Then, the CLIP model is utilized in the third stage to select the best-aligned caption for each image, filtering out inaccuracies. The process culminates with an LLM synthesizing captions from various perspectives into a comprehensive caption.
However, the captioning of rendered views (the combined second and third stages) for given 3D objects can falter with accidental views, producing captions that diverge significantly from the actual 3D object. In the worst-case scenarios, each rendering view might correspond to an incorrect object, leading to compounded errors when these captions are summarized by GPT4. One example is shown in Figure 3. Since GPT4 operates solely on text, it cannot correct these inaccuracies, resulting in captions riddled with hallucinated details.
Due to the versatility of 3D object geometries, determining which rendered views best reflect a 3D object’s characteristics is non-trivial. While measuring the geometric properties of 3D objects and computing their principal directions is feasible, positioning the camera orthogonally, as shown in the bottom-left example of Figure 2, is often suboptimal. Hence, we propose DiffuRank, which learns 3D priors from data to filter preferable rendered views by leveraging a pre-trained text-to-3D model. Our experiments demonstrate that DiffuRank efficiently enhances caption quality and reduces hallucinations with fewer renderings compared to using all available views.
@@ 3.2 DiffuRank Formulation
DiffuRank leverages a pre-trained text-to-3D diffusion model to rank rendered views based on their alignment with captions and the corresponding 3D information.
Formally, given a 3D object , a set of candidate captions , and a pre-trained diffusion model , the training objective of this diffusion model is to predict the 3D object conditioned on a text description . This corresponds to modeling the score function of the data distribution . Specifically, the diffusion model aims to minimize the objective:
where the noised input is defined as , for timestamp , randomly sampled Gaussian noise , and noise scheduling parameters defined as in [19]. Our intuition here is straightforward: a caption closely aligned with the 3D object’s characteristics (e.g., structure, colors, textures) should assist the diffusion model in accurately predicting from the same noised input , thus resulting in a lower score matching loss. By sampling multiple sets of for each caption , we measure the alignment score between the 3D object and captions as the average loss over these samples.
Initially, we generate candidate captions for by rendering it into multiple views and generating captions using an image captioning model . This captioning procedure aims to maximize the joint likelihood of the generated captions and the rendered image . Consequently, we estimate the alignment between the 3D object and all captions of a given rendering as , which is proportional to the alignment between the 3D object and the rendering itself:
We summarize the complete pipeline in Algorithm 1.
Specifically, we adopted shap-E as the text-to-3D diffusion model in our paper, and the above should be , where is the encoder (transmitter in [21]) to extract feature embeddings from given 3D object.
Furthermore, DiffuRank’s application is not confined to 3D captioning; because it is a general framework for measuring the alignment between two modalities received and output by a diffusion model. It can be seamlessly extended to other domains, such as 2D images. In section 5.3, we show an example where we apply DiffuRank to perform 2D VQA and beat CLIP model [49].
@@ 3.3 New 3D Captioning Framework
With the proposed DiffuRank, we establish a new 3D captioning pipeline, as shown in Figure 3. For given 3D object, we render it into 28 images, which are then captioned into 5 descriptions using an image-based captioning model. Following captioning, DiffuRank ranks the rendered views using a pre-trained text-to-3D model. This ranking enables the selection of the Top-6 rendered views for processing by a vision-language model, resulting in holistic captions that describe structure, form, color, and texture, with enhanced accuracy and detail.
To elaborate, our methodology integrates two distinct rendering strategies, as illustrated in Figure 4. The first strategy, derived from Cap3D [36], renders objects into 8 views against a uniform grey background, arranged horizontally around the object’s default orientation, with Blender ray-tracing render engine ‘CYCLES’. Concurrently, we apply a second technique from Shap·E [21], where 20 views are generated through randomized sampling after object normalization, set against a transparent background, with Blender real-time engine ‘EEVEE’. These 20 views, created following the Shap·E methodology, are instrumental in forming Shap·E latent codes, i.e. in Section 3.2. Altogether, this approach results in 28 distinct views for each object. Additionally, as grey and transparent backgrounds may accentuate or obscure details variably across objects, we observed that DiffuRank adeptly selects the views with the proper background that most effectively highlight object features, without manual intervention. Some examples are included in Appendix B.
Following this, the captioning model, BLIP2 [25], is employed to generate five captions for each view. These captions, alongside the pre-trained text-to-3D diffusion model, Shap·E [21], and the previously derived 3D latent code , undergo analysis in the DiffuRank process, as detailed in Algorithm 1. Subsequent to DiffuRank, the six views that demonstrate the highest alignment scores are chosen to input into GPT4-Vision for caption generation.